OpenStudy (anonymous):

write the answer in simplest form 16-6/10

OpenStudy (anonymous):

help

OpenStudy (igreen):

16 - 6 = ? @BLUEBELLA7887

OpenStudy (anonymous):

77/5 or 15.4

OpenStudy (igreen):

\(\dfrac{16-6}{10}\) \(\dfrac{10}{10}\) \(1\)

OpenStudy (igreen):

@jo.harvey How did you get that?

OpenStudy (anonymous):

thanks @jo.harvey

OpenStudy (anonymous):

16-0.6 = 15.4 than I made a fraction again

OpenStudy (igreen):

Oh wait.. Is it \(\dfrac{16-6}{10}\) or \(16-\dfrac{6}{10}\) @BLUEBELLA7887

OpenStudy (anonymous):

it's the second one :)

OpenStudy (igreen):

Oh, okay I was confused..lol.

OpenStudy (anonymous):

k thanks for the help